According to the rule of 70, if a country's real GDP per capita grows at an annual rate of 2% instead of 3%, it will take for that country to double its level of real GDP per capita. 

  • 30 additional years 

  • 11.7 additional years 

  • 35 fewer years 

  • 30 fewer years 

  • 35 additional years 

  • 23.3 additional years 

  • 23.3 fewer years 

  • 11.7 fewer years
Answer #1

According to the rule of 70, the expected time in which the real GDP per capita will double can be calculated by dividing the rate of growth by 70.

When the growth rate was 3% then it takes 70/3 = 23.3 years for the real GDP per capita to double.

When the growth rate falls to 2% then it takes 70/2 = 35 years for the real GDP per capita to double.

So, it takes 35 - 23.3 = 11.7 additional years for the real GDP per capita to double. Therefore, the correct answer is 'Option B'.
Answer #2

SOLUTION :


As per rule of 70.


Years required to double the amount = 70 / growth rate in % 


Hence,


Years required to double the GDP at 3% growth rate = 70 / 3 = 23.3 years


Years required to double the GDP at 2% growth rate = 70 / 2 = 35 years


So, additional years required to double the GDP at 5% growth rate instead of 7% :

= 35 - 23.3

= 11.7 additional years (ANSWER).

answered by: Tulsiram Garg

> Please correct, "So, additional years required to double the GDP at 5% growth rate instead of 7%" as "So, additional years required to double the GDP at 2% growth rate instead of 3%" :

Tulsiram Garg Thu, Oct 28, 2021 2:54 AM